OpenStudy (anonymous):

pls find the solution set of 3|x-5|=|2x-10|

OpenStudy (anonymous):

First take x = positive.. 3(x - 5) = (2x - 10) 3x - 15 = 2x - 10 x = 5 Now take x = negative ie -x 3(-x - 5) = 2(-x) - 10 -3x - 15 = -2x - 10 -x = 5 x = -5 is the answer...

OpenStudy (anonymous):

why isn't it 3(-x - 5) = -(2x-10) ?

OpenStudy (anonymous):

I just did not consider that one.. Ya it can be...

OpenStudy (anonymous):

By solving them what are the values you are getting??

OpenStudy (anonymous):

-5 cant be it... Try it